I need help with this question please:
Solve the system of equations
x + y = 5
x - y = -9

Found 2 solutions by Math_Teacher_37, nerdybill:
Answer by Math_Teacher_37(4)

We are trying to solve the following system of equations:
x + y = 5
x - y = -9
Notice that the "y" terms are opposite in sign so we can simply add the equations together, like this:
2x + 0y = -4
2x = -4
Divide both sides by 2: x = -4/2 = -2
Now, x = -2 so substitute into either equation. Since the first equation involves addition and it looks easier, let's go with this one. We have:
x + y = 5
-2 + y = 5
Add 2 to both sides:
-2 + 2 + y = 5 + 2
y = 7
The solution is: (-2, 7)

Answer by nerdybill(7384)

x + y = 5
x - y = -9
solve using the "addition" method:
x + y = 5
x - y = -9
--------------
2x +0y = -4
2x = -4
x = -2
.
substitute above into equation 1:
x + y = 5
-2 + y = 5
y = 7
.
solution:
x = -2
y = 7